OpenStudy (anonymous):

Holly is taking a multiple choice reading test. She has decided to guess on questions 3 and 7. On each question, there are four worded choices plus an "all of these" choice and a "none of these" choice. What is the probability that she answers question 3 correctly and question 7 incorrectly?

OpenStudy (anonymous):

I seeee.. there are 6 choices ..of which only 1 is correct .. Hence, probability of getting any question wrong is 5/6 ..so probability of getting question 3 wrong will be 5/6 .. and the same for question 7 .. Therefore probability of answering both question 3 and 7 incorrectly is (5/6) multiplied by (5/6) which equals .. 25/36

OpenStudy (kropot72):

The probability that she answers question 3 correctly and question 7 incorrectly is given by: \[\frac{1}{6}\times\frac{5}{6}=you\ can\ calculate\]